With a constant acceleration of magnitude 1.6 m/s2 due west , how long will it take a bus to travel 2.0 x 10^2 m , if the bus starts from rest?
Answer #1

a = 1.6m/s , u = 0m/s , t = ?

S = 200m

By using formula , S = ut + 1/2 at^2

=) 200 = 0 + 0.5 × 1.6 × t^2

=) t = √(200/0.8) = 15.8s